# Oriented Object Detection in Optical Remote Sensing Images using Deep Learning: A Survey

### v4

@@ 3 Detection Frameworks
It is widely accepted that object detection methods generally can be categorized into two groups: two-stage and one-stage detection (Liu et al., 2020; Zou et al., 2023). The former works in a coarse-to-fine paradigm while the latter completes classification and regression in one step thereby enjoying high efficiency but performing poorly on accuracy. Compared with the above two categories using anchor mechanisms, anchor-free methods directly detect objects without preset anchors. In addition, a series of DETR-based methods have merged recently, which view the detection process as a set prediction task, effectively eliminating several hand-craft components, e.g., NMS and anchor mechanism. Considering that each category has pros and cons, we divide the representative oriented object detectors into four categories: two-stage, one-stage, anchor-free, and DETR-based. Next, we will briefly review how each category addresses feature misalignment and spatial misalignment via deliberate framework design.
@@ 3.1 Two-Stage
Among numerous outstanding two-stage detectors (Ren et al., 2017; Lin et al., 2017; Cai and Vasconcelos, 2018; He et al., 2020; Qiao et al., 2021), Faster RCNN (Ren et al., 2017) armed with FPN (Lin et al., 2017) commonly serves as a benchmark due to its high accuracy and efficient design. As illustrated in Figure 3, it consists of the following pipeline: Feature Extraction, Region proposal networks (RPN), and Regions with CNN features (RCNN). In the first stage, a sparse set of high-quality region proposals, that can potentially contain objects, are generated via RPN (Chavali et al., 2016; Hosang et al., 2016). During the second stage, the region features are extracted for each proposal and then used for classification and refined regression using RCNN. Finally, several post-processing operations, such as NMS, are leveraged to output detection results (omitted in Figure 3).
@@
Its oriented version, termed as Rotated Faster RCNN or Faster RCNN OBB, predicts the orientation of each object by adding an extra channel in the regression branch. However, the naïve RPN only generates horizontal region proposals as regions of interest (RoIs), as shown in Figure 4(a). Apart from the feature misalignment caused by axis-aligned convolutions, another factor that may impair the final performance is the feature misalignment between horizontal RoI (HRoI) and OBB. Thus, the feature misalignment harms the feature representation significantly, making the detectors struggle to identify objects and regress precise OBBs yet inspiring successive innovations.
To cope with the feature misalignment, a variety of efforts are dedicated to generating rotated proposals and then adopting rotated RoI (RRoI) operators to extract spatial-aligned features. RRPN (Ma et al., 2018; Yang et al., 2018; Zhang et al., 2018) are designed with rotated anchors to fit the objects with different orientations. In addition to scales and aspect ratios, different orientation parameters are added to further generate additional rotated anchors, as shown in Figure 4(b). Such a scheme can alleviate the spatial misalignment thereby achieving better performance in terms of recall, but the redundant rotated anchors bring about expensive computation and memory consumption. To reduce the number of rotated anchors, RoI Transformer (Ding et al., 2019) retains the naïve RPN structure and introduces a lightweight learnable module, named RoI Learner. As shown in Figure 4(c), RRoI Learner can convert HRoIs to RRoIs directly. Such a design generates precise RRoIs without enormous rotated anchors, resulting in higher efficiency and accuracy. However, the RRoI Learner involves an additional stage including RoI operator and regression, making the networks complex and inefficient. As a result, Xie et al. (2021) design a simpler structure, named oriented RPN, to generate high-quality RRoIs from horizontal anchors directly (Figure 4(d)). Oriented RPN only contains a convolutional layer and two sibling convolutional layers. This lightweight module benefits from the proposed midpoint offset representation. For an oriented object, its midpoint offset representation consists of the corresponding external HBB and the offsets of vertexes w.r.t the midpoints of the external HBB. The external HBB can provide bounded constraint for OBB and the offsets can avoid the angle regression. In addition, midpoint offset representation retains the horizontal regression mechanism, making the training more stable than regressing OBBs from horizontal anchors. Benefiting from the design of oriented RPN and midpoint offset representation, Oriented RCNN can achieve competitive accuracy to advanced two-stage detectors and reach approximate efficiency to one-stage detectors.
@@ 3.2 One-Stage
Different from two-stage detection frameworks working in a coarse-to-fine paradigm, one-stage detectors complete classification and regression in one step Lin et al. (2020). As illustrated in Figure 5, one-stage detectors first extract multi-level feature maps and then predict the class probabilities and locations for each anchor per spatial location. Since the removal of RPN and RoI operators, one-stage detectors exhibit more severe feature misalignment compared to two-stage ones. Thus, a series of one-stage algorithms are developed to alleviate the dilemma, such as RDet (Yang et al., 2021b), ADT-Det (Zheng et al., 2021), and SA-Net (Han et al., 2022).
RDet (Yang et al., 2021b) adopts a feature refinement module (FRM) to align features. First, RDet transforms the horizontal anchors into rotated anchors, which can provide more accurate positional and oriented information. Then, FRM employs pixel-wise feature interpolation to integrate features from five locations (i.e., one center and four corners) of the corresponding refined rotated anchors. Similarly, SA-Net (Han et al., 2022) uses alignment convolution (AlignConv) to align the features. As a variant of deformable convolution (Dai et al., 2017), the offset field of AlignConv is inferred from the guide of rotated anchors. These schemes work in a coarse-to-fine paradigm to align features but are noticeably different from the RRoI operator. The major difference lies in that the FRM or AlignConv follows a full convolution structure and has fewer sampling points than the RRoI operator, making it more efficient.
@@ 3.3 Anchor-Free
The above two categories follow the anchor paradigm which suffers from spatial misalignment between horizontal anchor and OBB. To tackle the above issues, a constellation of anchor-free methods is developed to find objects without preset anchors. These methods eliminate anchor-related hyper-parameters, showing potential in the generalization to wide applications (Zhang et al., 2020b). According to the representation of OBB, anchor-free methods can be divided into keypoint-based methods (Guo et al., 2021; Wei et al., 2020) and center-based methods (Guan et al., 2021; He et al., 2021; Lin et al., 2019; Xiao et al., 2020; Yi et al., 2021; Zhang et al., 2022b; Zhou et al., 2020; Zhao et al., 2021).
@@
The keypoint-based methods first locate a set of adaptive or self-constrained key points and then circumscribe the spatial extent of the object, as shown in Figure 6(a). O-DNet (Wei et al., 2020) first locates the midpoints of four sides of the OBB by regressing the offsets from the center point. Then, two sets of opposite midpoints are connected to form two mutually perpendicular midlines which can be decoded to get the representation of OBB. In addition, a self-supervision loss is designed to constrain the perpendicular relationship between two middle lines and a collinear relationship between the center point and two opposite midpoints. Following the RepPoints (Yang et al., 2019b), CFA (Guo et al., 2021) utilizes the deformable convolution (Dai et al., 2017) to generate a convex hull for each oriented object. The convex hull is represented by a set of irregular sample points bounding the spatial extent of an object, which are determined by the designed Convex Intersection over Union (CIoU) loss. To alleviate feature aliasing between densely packed objects, convex-hull set splitting and feature anti-aliasing strategies are designed to refine the convex-hulls and adaptively optimal feature assignment. Furthermore, to predict the high-quality oriented reppoints, Oriented RepPoints (Li et al., 2022b) further designs an Adaptive Points Assessment and Assignment (APAA) scheme to measure the quality of reppoints. Such a scheme assesses each set of reppoints from four aspects, including classification, localization, orientation alignment, and point-wise correlation. As a result, the high-quality reppoints obtained via APAA enable Oriented RepPoints to achieve state-of-the-art performance among anchor-free methods.
@@
The center-based methods generally generate multiple probabilistic heatmaps and a series of feature maps. As shown in Figure 6(b), the heatmaps provide a set of candidates (peak points) as coarse center points. Meanwhile, the feature maps regress transformation parameters to represent the OBB. Nowadays, most center-based methods are dedicated to designing a variety of OBB representations to address PoA problems (Guan et al., 2021; He et al., 2021; Lin et al., 2019; Xiao et al., 2020; Yi et al., 2021; Zhang et al., 2022b; Zhou et al., 2020; Zhao et al., 2021). However, the center-based methods typically follow the one-stage paradigms and tend to predict coarse locations due to feature misalignment, while state-of-the-art methods generally contain one or multiple refined stages to improve performance. Thus, an effective scheme for performance improvement is leveraging anchor-free methods to generate coarse proposals that are then refined to generate high-quality proposals or detection results, e.g., AOPG (Cheng et al., 2022a), DEA (Liang et al., 2022). This could be attributed to the fact that the anchor-free rotated proposal generation scheme can not only generate accurate proposals but also avoid the spatial misalignment caused by horizontal anchors. Nevertheless, there is still a significant performance gap between plain center-based oriented methods and other state-of-the-art methods, necessitating further research.
@@ 3.4 DETR-Based
In addition to the above convolution-based methods, DETR-based detectors show great potential and achieve state-of-art performance in the detection community, including DETR (Carion et al., 2020) and its variants (Zhu et al., 2021; Sun et al., 2021; Gao et al., 2021). Based on DETR (Carion et al., 2020), ODETR (Ma et al., 2021) was proposed to utilize the Transformer for the oriented object detection task. In addition, the depthwise separable convolutions (Sifre and Mallat, 2013; Chollet, 2017; Haase and Amthor, 2020) is introduced to replace the computationally complex self-attention mechanism, making networks more lightweight and speeding up the training. To tackle feature misalignment, Dai et al. (2022) propose AO2-DETR by improving Deformable DETR (Zhu et al., 2021), which design an oriented proposal generation mechanism and an adaptive oriented proposal refinement (OPR) module for aligning the features. Recently, several improved DETR-based detectors have been proposed for generic object detection, e.g., DN-DETR (Li et al., 2022a), DAB-DETR (Liu et al., 2022a), and DINO (Zhang et al., 2022), bringing about dramatic breakthroughs in accuracy and convergence speed. Zeng et al. (2024) propose ARS-DETR that attempts to exploit DINO (Zhang et al., 2022) to oriented object detection tasks. Compared to other advanced oriented object detectors, it achieves greater detection accuracy in the more rigorous metric (i.e., AP), but lags behind in the standard metric (i.e., AP). Worse still, the long training convergence time and heavy computation cost are still open problems.
@@ 3.5 Discussion
Feature misalignment and spatial misalignment will seriously impair the performance of oriented object detection. To tackle these issues, an enormous amount of research effort has contributed to modifying the detection frameworks. Existing two-stage detectors typically design efficient and precisely oriented proposal generation modules, which can extract rotated aligned features with RRoI operators. Similarly, one-stage detectors are inclined to add an extra refined stage for feature alignment. Thus, the above two schemes can empower the detectors to mine the rotation-related information, thereby enhancing the semantic representation of oriented objects. Nevertheless, the modules are generally troubled by spatial misalignment. To alleviate this issue, anchor-free detectors discard the anchor mechanism. However, current anchor-free methods generally show limited performance due to feature misalignment. In addition, DETR-based methods provide a new detection paradigm and receive widespread concerns. However, the exploration of DETR-based methods in oriented object detection is not comprehensive enough, requiring further research to accelerate training convergence and reduce computation overheads.
Well-designed detection frameworks are conducive to alleviating feature and spatial misalignment but fail to address the PoA problems. Besides, the extracted features are not equipped with rotation-invariance as the convolution operators are axis-aligned. To cope with the above dilemma, the suitable OBB regression and the powerful feature representation of oriented objects are also widely studied, since they can be seamlessly integrated into various detection frameworks. Next, we will discuss the OBB regression and feature representation.
@@ 4 OBB Regression
Oriented object detectors typically locate objects in a regression fashion. The customized regression head predicts the orientation parameters when using the most frequently used -based representation. Unfortunately, such a regression paradigm suffers from several limitations, including inconsistency between metric and loss, and angular boundary discontinuity. For more details about the above issues please refer to Section A of the Appendix or corresponding papers (Qian et al., 2021, 2022; Yang et al., 2021c, 2022b; Xu et al., 2021a). To tackle these issues, existing oriented object detection methods usually develop new loss functions or redesign OBB representation. Next, we will briefly introduce these approaches and discuss their pros and cons.
@@ 4.1 Regression Loss
(1) Inconsistency between Metric and Loss. The inconsistency between metric and loss generally implies that an optimum choice for the regression task may not guarantee a high localization accuracy in terms of IoU. To address this issue, existing generic object detectors generally introduce IoU-induced loss functions, such as GIoU (Rezatofighi et al., 2019) and DIoU (Zheng et al., 2020). However, these IoU-induced losses cannot be incorporated directly into oriented object detection due to the in-differentiable nature of RIoU (Yang et al., 2021c). Thus, several differentiable functions are designed to approximate RIoU loss (Yang et al., 2021b, 2019a, 2022a). Chen et al. (2020) proposed PIoU by introducing a differentiable kernel function, which accumulates the contribution of interior overlapping pixels to approximate the intersection area. Several solutions (Yang et al., 2019a, 2022a, 2021b) integrated RIoU as a loss weight of the regression loss:
denotes the commonly used smooth L1 loss (Ren et al., 2017). is a loss function related to RIoU, e.g., . Such a loss can be divided into two parts, a normalized regression loss controlling the direction of gradient propagation, and a scalar adjusting gradient magnitude. When the RIoU is close to 1, , and is approximately equal to 0, which can effectively alleviate the inconsistency between the metric and regression loss. Apart from designing regression loss to approach RIoU, Ming et al. (2024) analyze the gradient and propose a Gradient Calibration Loss (GCL). GCL build a corrected gradient w.r.t RIoU, angular error, and scale, and then calculate the optimized regression loss through integration. However, angle regression still suffers from the problem of PoA.
@@
Modulated rotated loss (Qian et al., 2021, 2022). The modulated Rotated loss adds an extra loss item based on the naïve regression loss to eliminate the angular boundary discontinuity. Specifically, it first transforms the original predicted OBB to another form 111It is noteworthy that the modulated rotated loss is customized for OBB representation under the OpenCV definition, which is intractable to the exchangeability of edges and PoA., and then take the minimum of their regression loss, i.e., , where denotes the corresponding ground-truth OBB. Such a scheme can adaptively choose the appropriate representation of the predicted OBB making the smallest loss value, thereby effectively eliminating the sudden increase in loss when the angle approaches its boundary and ensuring the loss continuous. However, the gap between metric and loss still exists.
Angle coder (Yang and Yan, 2020; Yang et al., 2021a; Yu and Da, 2023). Yang and Yan (2020) proposed a new baseline by transforming the angular regression task into a classification problem. The angle is discretized into a certain number of intervals, and then a discrete angle is predicted by classification. Besides, to increase the error tolerance to adjacent angles and handle the PoA, the circular smooth label (CSL) technique is designed. Although CSL eliminates the boundary discontinuity, its heavy prediction layer harms the efficiency and the square-like problem remains unsolved. To tackle these issues, Yang et al. (2021a) further adopted Densely Coded Labels (DCL) by introducing a lightweight prediction layer, which reduces the code length. Moreover, Angle Distance and Aspect Ratio Sensitive Weighting (ADARSW) were designed to improve the accuracy of square-like objects. However, the hyper-parameters have a significant impact on the performance of CSL and DCL. Even worse, the optimal settings on different datasets are also different thereby requiring laborious tuning. To solve this problem, Yu and Da (2023) designed a differentiable angle coder, named Phase-Shifting Coder (PSC). PSC encodes the angle into a periodic phase to solve the boundary discontinuity problem. Moreover, to further solve the square-like problem, PSCD was proposed by mapping the angle into the phases of different frequencies, which is the advanced dual-frequency version of PSC.
@@
Gaussian distribution based methods (Yang et al., 2021c, d, 2022, 2022b). The Gaussian distribution based methods provide a unified and elegant solution to the problems of boundary discontinuity and the square-like problem. First, the OBB is converted to a 2-D Gaussian distribution , where and is a matrix associated with :
@@
Then a distance function is used to measure two Gaussian distributions, such as Gaussian Wasserstein Distance (GWD) (Yang et al., 2021c) or Kullback-Leibler Divergence (KLD) (Yang et al., 2021d). Furthermore, the measure between Gaussian distributions is converted into an approximate IoU loss using a nonlinear transformation to obtain consistency between the metric and the regression loss. The merit of Gaussian distribution is that the angle is encoded by trigonometric function thereby not constrained by PoA. Moreover, the OBB parameters are joint-optimized dynamically so that they can influence each other during training. However, both GWD and KLD only maintain value-level consistency instead of trend-level consistency between RIoU and regression loss. To ameliorate this problem, Yang et al. (2022) design the KFIoU loss based on the Kalman filter to achieve the best trend-level alignment with RIoU, which is differentiable and does not require additional hyperparameters. Based on the mechanism of RIoU, KFIoU calculates the overlapping area between two Gaussian distributions via a Kalman filter. Benefiting from the trend-level alignment strategy, KFIoU achieves better performance than GWD and KLD.

### v5

@@ 3 Detection Frameworks
It is widely accepted that object detection methods can be categorized into two primary groups: two-stage and one-stage detection [96, 205]. The former works in a coarse-to-fine paradigm, whereas the latter accomplishes classification and regression in one step, thereby exhibiting high efficiency but performing poorly on accuracy. In contrast to the aforementioned two categories that rely on anchor mechanisms, anchor-free methods directly detect objects without the need for predefined anchors. In addition, a series of DETR-based methods have merged recently, regarding the detection process as a set prediction task, thereby effectively eliminating several hand-craft components, e.g., NMS and anchor mechanism. Considering that each category has pros and cons, we divide the representative oriented object detectors into four categories: two-stage, one-stage, anchor-free, and DETR-based. Several milestone methods are presented in Fig. 6. Next, we will concisely review how each category addresses feature misalignment and spatial misalignment via deliberate framework design.
@@ 3.1 Two-Stage
Among numerous prominent two-stage detectors [130, 131, 77, 17, 49, 122], Faster RCNN [130, 131] armed with FPN [77] commonly serves as a benchmark due to its exceptional accuracy and efficient design. As depicted in Fig. 7a, its workflow consists of the following pipeline: Feature Extraction, Region proposal networks (RPN), and Regions with CNN features (RCNN). In the first stage, a sparse set of high-quality region proposals, that can potentially contain objects, are generated via RPN [5, 44]. During the second stage, the region features are extracted for each proposal and then used for classification and refined regression via RCNN. Finally, several post-processing operations, such as NMS, are leveraged to finalize the detection results (omitted in Fig. 7a). Its oriented version, termed as Rotated Faster RCNN or Faster RCNN OBB, predicts the orientation of each object by adding an extra channel in the regression branch.
@@
However, the naïve RPN only generates horizontal region proposals as regions of interest (RoIs), as shown in Fig. 9a. Apart from the feature misalignment caused by axis-aligned convolutions, another factor that may impair the final performance is the feature misalignment between horizontal RoI (HRoI) and OBB, as shown in Fig. 8a. The feature misalignment significantly harms feature representation, making the detectors struggle to identify objects and regress precise OBBs yet inspiring successive innovations.
To cope with the feature misalignment, various efforts are dedicated to generating rotated proposals and then adopting rotated RoI (RRoI) operators to extract spatial-aligned features, as shown in Fig. 8b. RRPN [113, 190, 210] incorporates rotated anchors to accommodate objects of various orientations. In addition to scales and aspect ratios, different orientation parameters are added to further generate additional rotated anchors, as shown in Fig. 9b. Such a scheme can alleviate the spatial misalignment (shown in Fig. 4c) thereby achieving better performance in terms of recall. However, the redundant rotated anchors bring about expensive computation and memory consumption.
To reduce the number of rotated anchors, RoI Transformer [29] retains the naïve RPN structure to alleviate spatial misalignment (shown in Fig. 4b), and then introduces a lightweight RoI Learner module. As shown in Fig. 9c, RRoI Learner converts HRoIs directly into RRoIs, generating precise RRoIs without enormous rotated anchors, thus enhancing efficiency and accuracy. Yet, the added complexity of the RoI Learner, including an extra RoI operator and regression stage, makes the network less efficient.
Consequently, [174] design a simpler structure, Oriented RCNN, to generate high-quality RRoIs from horizontal anchors directly, as shown in Fig. 9d. This lightweight module benefits from the proposed Midpoint Offset representation, which includes the corresponding external HBB and the offsets of vertexes w.r.t the midpoints of the external HBB. This representation maintains horizontal regression mechanisms, ensuring stable training compared to regressing OBBs from horizontal anchors. Benefiting from the design of oriented RPN and Midpoint Offset representation, Oriented RCNN can achieve competitive accuracy to advanced two-stage detectors and reach approximate efficiency to one-stage detectors.
The excellent design of RoI Transformer and Oriented RCNN tackles the spatial misalignment and feature misalignment in two-stage detectors, laying the foundation for subsequent research in this field. Numerous subsequent two-stage detectors adopt RoI Transformer or Oriented RCNN as their baseline framework, harnessing exceptional feature representation (e.g., ReDet [48], RVSA [169], ARC [121], STD [191]) or crafting superior OBB representation (e.g., Gliding Vertex [176], RPGAOD [123], QPDet [183]) and loss functions (e.g., FRIoU [124], CGCDet [170], GCL [111]) to enhance their performance. These methods will be discussed in detail in the subsequent section.
@@ 3.2 One-Stage
As illustrated in Fig. 7b, one-stage detectors first extract multi-level feature maps and then predict the class probabilities and locations for each anchor per spatial location. Due to the absence of RPN and RoI operators, one-stage detectors encounter more severe feature misalignment than two-stage ones. Thus, a series of one-stage algorithms are developed to alleviate the dilemma, such as R3Det [193] and S2A-Net [47].
R3Det [193] adopts a feature refinement module (FRM) to align features. First, R3Det transforms the horizontal anchors into rotated anchors, which can provide more accurate positional and oriented information. Then, FRM employs pixel-wise feature interpolation to integrate features from five locations (i.e., one center and four corners) of the corresponding refined rotated anchors, as shown in Fig. 3b. Similarly, S2A-Net [47] aligns features using Alignment Convolution (AlignConv), a variant of deformable convolution [26]. The offset field of AlignConv is inferred from the guidance of rotated anchors. Both FRM and AlignConv operate in a coarse-to-fine manner but differ significantly from the RRoI operator. Notably, they follow a full convolution structure with fewer sampling points, resulting in increased efficiency.
Based on S2A-Net, CFL [144] introduces a Spatial Transform Selection (STS) strategy and a Critical Feature Sampling (CFS) module. STS dynamically assigns labels by calculating IoU thresholds based on aspect ratios, angle differences, and the initial IoU threshold determined by ATSS [206]. The adaptable IoU threshold controls the number of samples assigned to the easy object while ensuring enough positive samples for hard one with large aspect ratio and angle difference. CFS incorporates a deformable convolution in which the sampling positions are derived from initial detection results (center point, vertices, and midpoints) combined with a learnable offset field.
Both R3Det and S2A-Net utilize the extra refined head of for feature alignment, making them popular baseline choices for one-stage detectors. These methods further solve spatial misalignment and OBB regression problem via innovative sample assignment strategies (e.g., DAL [115], DCFL [175]) or regression loss (e.g., DCL [187], FoRDet [236], GWD [198], KLD [200], PSC [185, 186]). In the following section, we will provide a detailed introduction to these methods.
@@ 3.3 Anchor-Free
The above two categories follow the anchor paradigm which suffers from spatial misalignment between horizontal anchor and OBB. To tackle the above issues, a constellation of anchor-free methods are developed to detect objects without relying on preset anchors. These methods eliminate anchor-related hyper-parameters, showing potential in the generalization to wide applications [206]. According to the representation of OBB, anchor-free methods can be divided into keypoint-based and center-based methods.
@@
Keypoint-based methods first locate a set of adaptive or self-constrained key points and then circumscribe object’s spatial extent, as shown in Fig. 10a. For instance, O2-DNet [166] first locates the midpoints of four sides of the OBB by regressing the offsets from the center point. Then, two sets of opposite midpoints are connected to form two mutually perpendicular midlines, which are decoded to represent the OBB. In addition, a self-supervision loss constrains the perpendicular relationship between two middle lines and a collinear relationship between the center point and two opposite midpoints. Following the RepPoints [188], CFA [38, 40] utilizes the deformable convolution [26] to generate a convex hull for each oriented object. The convex hull, represented by irregular sample points, is refined using a Convex Intersection over Union (CIoU) loss. To alleviate feature aliasing between densely packed objects, convex-hull set splitting and feature anti-aliasing strategies are designed to refine the convex-hulls and adaptively optimal feature assignment.
Furthermore, to predict the high-quality oriented reppoints, Oriented RepPoints [74] designs an Adaptive Points Assessment and Assignment (APAA) scheme to measure the quality of reppoints. APAA assesses reppoints across four dimensions (classification, localization, orientation alignment, and point-wise correlation) to select high-quality ones, without adding computational burden during inference. Subsequently, [197] propose a dynamic information aggregation (DIA) module based on multi-head self-attention mechanism [151]. By mining the relationships between reppoints, DIA not only helps to obtain more accurate positions of reppoints but also enriches the feature representations, thereby further boosting localization accuracy.
@@
Center-based methods generally generate multiple probabilistic heatmaps and a series of feature maps. As shown in Fig. 10b, the heatmaps provide a set of candidates (peak points) as coarse center points, while the feature maps regress transformation parameters to represent the OBB. Nowadays, most center-based methods are dedicated to designing a variety of OBB representations to address the PoA problems, including CHPDet [221], GGHL [57], DHRec [116]. However, these methods typically follow the one-stage paradigms and tend to predict coarse locations due to feature misalignment, while state-of-the-art methods generally contain one or multiple refined stages to improve performance.
Hence, an effective scheme for boost performance is leveraging anchor-free methods to generate coarse detection results that are then refined via subsequent feature alignment stage, e.g., AOPG [19], DEA [80], DRDet [216], TS-Conv [55]. AOPG [19] initially produces coarse oriented boxes via the rotated FCOS manner [146], subsequently refining them into high-quality oriented proposals. DEA [80] leverages two parallel branches, which generate proposals using anchor-free and anchor-based approaches respectively, followed by an interactive sample screening procedure to select high-quality training samples.
In contrast to the above two methods that capitalize on the merits of anchor-free techniques to mitigate spatial misalignment and facilitate appropriate sample assignment, DRDet [216] and TS-Conv [55] concentrate on feature refinement. DRDet [216] adopts two perpendicular rotated lines to represent OBB. Then, an orientation-guided feature encoder (OFD) is designed to encode the orientation-aware information into refined features along each rotated line. Compared to rectangular feature, the line features extracted from OFD can introduce less noise and alleviate the feature aliasing caused by overlapping objects. TS-Conv [55] designs different sampling offsets for localization and classification to alleviate the task misalignment problem (i.e., localization and classification tasks may focus on different feature regions [138]). The sampling offsets are restricted by initial OBBs predicted via an anchor-free manner, allowing for dynamic adaptation to objects with various shapes.
@@ 3.4 DETR-Based
In addition to the above convolution-based methods, DETR-based detectors show great potential and achieve state-of-the-art performance in the detection community, including DETR [13] and its variants [218, 135, 41]. Based on DETR [13], O2DETR [110] is proposed to utilize the Transformer for the oriented object detection task. In addition, the depthwise separable convolutions [139, 8, 43] is introduced to replace the computationally complex self-attention mechanism, making networks more lightweight and speeding up the training. To tackle feature misalignment, [25] propose AO2-DETR by improving Deformable DETR [218], which design an oriented proposal generation mechanism and an adaptive oriented proposal refinement (OPR) module for aligning the features. Recently, several improved DETR-based detectors have been proposed for generic object detection, e.g., DN-DETR [104], DAB-DETR [91], and DINO [213], bringing about dramatic breakthroughs in accuracy and convergence speed. ARS-DETR [207] attempts to exploit DINO [213] for oriented object detection tasks. Compared to other advanced oriented object detectors, it achieves greater detection accuracy in the more rigorous metric (i.e., AP75), but lags in the standard metric (i.e., AP50). Worse still, the long training convergence time and heavy computation cost are still open problems.
@@ 3.5 Discussion
Feature misalignment and spatial misalignment will seriously impair the performance of oriented object detection. To tackle these issues, substantial research has contributed to modifying the detection frameworks. Existing two-stage detectors typically design efficient and precisely oriented proposal generation modules, leveraging RRoI operators to extract rotated aligned features. Similarly, one-stage detectors are inclined to incorporate an extra refined stage for feature alignment. Thus, the above two schemes can empower the detectors to mine the rotation-related information, thereby enhancing the semantic representation of oriented objects.
Nevertheless, spatial misalignment remains a persistent issue. Anchor-free detectors address this by eliminating the anchor mechanism and, in advanced methods, incorporating extra feature refinement stages to further mitigate feature misalignment. In addition, DETR-based methods provide a new detection paradigm and receive widespread concerns. However, the exploration of DETR-based methods in oriented object detection is not comprehensive enough, requiring further research to accelerate training convergence and reduce computation overheads.
Well-designed detection frameworks are conducive to alleviating feature and spatial misalignment but fail to address the PoA problems. Besides, the extracted features are not equipped with rotation-invariance as the convolution operators are axis-aligned. To overcome these dilemmas, the suitable OBB regression and the powerful feature representation of oriented objects are also widely studied, since they can be seamlessly integrated into various detection frameworks. Next, we will discuss the OBB regression and feature representation.
@@ 4 OBB Regression
Oriented object detectors typically locate objects in a regression fashion. Specifically, the customized regression head predicts the orientation parameters when using the most frequently used -based representation. Unfortunately, such a regression paradigm suffers from several limitations, including inconsistency between metric and loss, and angular boundary discontinuity. For more details about the above issues please refer to Sec. A of the Appendix or corresponding papers [125, 126, 198, 202, 176]. To tackle these issues, existing oriented object detection methods usually develop novel loss functions or or alternative representations for OBB. Several representative methods are shown in Fig. 11. Next, we will briefly introduce them and discuss their pros and cons.
@@ 4.1 Regression Loss
(1) Inconsistency between Metric and Loss. The inconsistency between metric and loss generally implies that an optimum choice for the regression task may not guarantee a high localization accuracy in terms of IoU. To bridge this gap, existing generic object detectors generally introduce IoU-induced loss functions, such as GIoU [134] and DIoU [220]. However, these IoU-induced losses cannot be incorporated directly into oriented object detection due to the in-differentiable nature of RIoU [198]. Thus, several differentiable functions are designed to approximate RIoU loss [199, 193, 194]. PIoU [6] introduces a differentiable kernel function that accumulates the contribution of interior overlapping pixels to approximate the intersection area. Several solutions [199, 194, 193] integrate RIoU as a loss weight of the regression loss:
denotes the commonly used smooth L1 loss [130, 131]. is a loss function related to RIoU, e.g., . Such a loss is composed of a normalized regression loss controlling the direction of gradient propagation, and a scalar adjusting gradient magnitude. When the RIoU is close to 1, , and is approximately equal to 0, effectively mitigating the inconsistency between the metric and regression loss.
Apart from designing regression loss to approach RIoU, [111] analyze the gradient of RIoU loss and propose the Gradient Calibration Loss (GCL). GCL constructs a corrected gradient w.r.t RIoU, angular error, and scale, and then calculates the optimized regression loss through integration. Despite these efforts, angle regression still faces challenges, particularly the problem of PoA.
@@
Modulated rotated loss [125, 126]. The modulated Rotated loss adds an extra loss item based on the naïve regression loss to eliminate the angular boundary discontinuity. Specifically, it first transforms the original predicted OBB to another form 111It is noteworthy that the modulated rotated loss is customized for OBB representation under the OpenCV definition, which is intractable to the exchangeability of edges and PoA., and then take the minimum of their regression loss, i.e., , where denotes the corresponding GT. Such a scheme can adaptively choose the appropriate representation of the predicted OBB making the smallest loss value, thereby mitigating the sudden increase in loss near angular boundaries. However, this approach does not fully resolve the metric-loss inconsistency.
Angle coder [192, 187, 185, 186]. The circular smooth label (CSL) approach discretizes angles into intervals and predicts a discrete angle via classification [192]. Besides, to increase the error tolerance to adjacent angles and handle the PoA, CSL uses a window function for angle label smoothing. Although CSL eliminates the boundary discontinuity, its heavy prediction layer harms the efficiency. To tackle these issues, [187] further adopt Densely Coded Labels (DCL) to reduce the code length. Furthermore, [158] analyze the limitations of CSL when directly applying continuous Focal Loss functions [78, 79] to the soft labels of angle classification. Specifically, when the label , the extreme point of the derivative of Focal Loss functions is not at , causing inaccurate angle prediction. Thus, Gaussian focal-CSL (GF-CSL) is designed to obtain more accurate angle predictions with higher responses at peaks via adaptive Gaussian attenuation on the negative angle categories. However, the hyper-parameters have a significant impact on the performance of these methods. Even worse, the optimal settings on different datasets are also different thereby requiring laborious tuning.
To solve this problem, Yu and Da [185, 186] design a differentiable angle coder, named Phase-Shifting Coder (PSC). PSC encodes the angle into a periodic phase to solve the boundary discontinuity problem. Moreover, an advanced version, PSCD, maps angles into phases of different frequencies to further solve the square-like problem.
@@
Gaussian distribution based methods [198, 203, 200, 202]. The Gaussian distribution based methods provide a unified and elegant solution to the problems of boundary discontinuity and the square-like problem. First, the OBB is converted to a 2-D Gaussian distribution , where and is a matrix associated with :
@@
Then a distance function is used to measure two Gaussian distributions, such as Gaussian Wasserstein Distance (GWD) [198] or Kullback-Leibler Divergence (KLD) [200]. Furthermore, the measure is converted into an approximate IoU loss using a nonlinear transformation to obtain metric-loss consistency. The merit of Gaussian distribution is that the angle is encoded by trigonometric function thereby not constrained by PoA. Moreover, the OBB parameters are joint-optimized dynamically so that they can influence each other during training.
Despite their advantages, both GWD and KLD only maintain value-level consistency instead of trend-level consistency between RIoU and regression loss. To achieve better trend-level alignment, the KFIoU loss is proposed, which is differentiable and does not require additional hyperparameters. KFIoU can calculate the overlapping area between two Gaussian distributions, resulting in improved performance compared to GWD and KLD.